OpenStudy (anonymous):

Using complete sentences, explain how you would use the quadratic formula to solve x2 + 5x = –2. Why is the quadratic formula the best method to use?

OpenStudy (anonymous):

x2 + 5x +2=0 its in form ax2+bx+c=0 x=-b+root(D)/2a or -b-root(D)/2a D=b^2-4ac